# A Real-time Anomaly Detection Method for Robots based on a Flexible and Sparse Latent Space

### v2

@@ Abstract
The growing demand for robots to operate effectively in diverse environments necessitates the need for robust real-time anomaly detection techniques during robotic operations. However, deep learning-based models in robotics face significant challenges due to limited training data and highly noisy signal features. In this paper, we present Sparse Masked Autoregressive Flow-based Adversarial AutoEncoders model to address these problems. This approach integrates Masked Autoregressive Flow model into Adversarial AutoEncoders to construct a flexible latent space and utilize Sparse autoencoder to efficiently focus on important features, even in scenarios with limited feature space. Our experiments demonstrate that the proposed model achieves a 4.96% to 9.75% higher area under the receiver operating characteristic curve for pick-and-place robotic operations with randomly placed cans, compared to existing state-of-the-art methods. Notably, it showed up to 19.67% better performance in scenarios involving collisions with lightweight objects. Additionally, unlike the existing state-of-the-art model, our model performs inferences within 1 millisecond, ensuring real-time anomaly detection. These capabilities make our model highly applicable to machine learning-based robotic safety systems in dynamic environments. The code will be made publicly available after acceptance.
@@ 1 Introduction
Neural network-based time series anomaly detection approaches have been actively researched in recent times (Pang et al., 2021). Unsupervised learning, which uses only normal data while training the model, is predominantly used in this field (Pang et al., 2021) since the occurrence of anomalies is very rare and significant effort is required to identify and label data. Models such as Long Short-Term Memory-based Variational AutoEncoder (LSTM-VAE) (Park et al., 2018), MultiVariate Time-series Flow (MVT-Flow) (Brockmann et al., 2024), Anomaly transformer (Xu et al., 2022) are proposed to efficiently detect anomalies using unsupervised learning.
@@
Anomaly transformer achieves success in the field of anomaly detection by utilizing transformer encoders (Vaswani et al., 2017) to capture both temporal and global time-series features of the data. It can effectively detect anomalies because the transformer can capture global time-series information that recurrent neural networks are unable to retain. However, due to the algorithm’s time and space complexity of $O(L^{2}N)$, where $L$ represents the length of the time series sequence and $N$ denotes the number of features, it requires substantial computational time and memory space. This becomes a drawback when performing real-time anomaly detection with long time steps. In robotics, the high computational cost of the method is a critical weakness since it has to be implemented on the embedded computing unit for real-time anomaly detection. Specifically, using numerous sensor signals increases the computational load, rendering it unsuitable for use on robots, particularly low-cost commercial robots (Yen et al., 2019). Additionally, the model’s large size requires extensive datasets for training, but building these datasets through robot operation is challenging due to the time-intensive data collection process. (Soori et al., 2023).
@@
In this paper, we propose a real-time multivariate time series anomaly detection model, Sparse Masked Autoregressive Flows-based Adversarial AutoEncoder (Sparse MAF-AAE), designed to address the limitations of conventional anomaly detection models. First, we enhance the adaptability of the model’s latent space for datasets, designing it to have a flexible posterior and prior distributions. This flexibility is achieved by transforming the posterior and prior distributions within Adversarial AutoEncoder (AAE) (Makhzani et al., 2016) using Masked Autoregressive Flow (MAF) (Papamakarios et al., 2017), a type of NF. The combination of AAE and MAF enables the model to learn a latent space that captures intricate data structures, performing well even with a small amount of training data. Additionally, this approach enhances robustness against out-of-distribution data by integrating Masked Autoencoder for Distribution Estimation (MADE) (Germain et al., 2015) within MAF, utilizing a bottleneck architecture (Kirichenko et al., 2020). Second, to preserve essential signal characteristics when the number of signal features is small, we incorporate the concept of Sparse autoencoder (Ng et al., 2011) by creating a latent space larger than the input while applying constraints that limit parameter updates. This configuration enables logical compression of the input data and ensures effective feature preservation for constructing an appropriate latent space from sparse input sensor signals. As a result, it enhances the model’s effectiveness in detecting anomalies in robotic operations with limited sensor signals. Finally, we demonstrated that our methods are feasible for real-time multivariate anomaly detection in robots by constructing models with simple architecture, achieving suitable elapsed times for real-time inferences. To the best of our knowledge, this is the first work to integrate a flow-based model, AAE, and Sparse autoencoder within a unified framework for signal reconstruction-based anomaly detection.
@@
Comparative experiments are conducted to evaluate the real-time multivariate anomaly detection performance based on a modified version of the voraus-AD dataset (Brockmann et al., 2024). The dataset offers time series signals of the manipulator when a pick-and-place operation is performed with 12 anomaly types caused by heterogeneous factors. To implement the real-time detection, the sliding window is applied for the implemented anomaly detection methods. Results showed that our model achieved the highest area under the ROC curve (AUROC) score among the comparison models, while also maintaining an appropriate computational time for real-time multivariate time-series anomaly detection in robotic applications.
@@ 4.4 Results
Table 3: AUROC percentages of real-time anomaly detection models on 100 Hz torque signals trained with 948 (100%) samples.
Model
Anomaly
transformer
LSTM-VAE
MVT-Flow
(previous SOTA)
Sparse MAF-AAE
Add. Friction 69.63 81.60 ± 1.52 87.20 96.29 ± 0.03
Miscommutation 58.35 67.53 ± 1.75 90.96 75.87 ± 0.11
Misgrip can 62.53 97.99 ± 2.28 99.74 99.82 ± 0.02
Losing can 56.83 88.48 ± 1.35 88.20 89.03 ± 0.11
Add. axis weight 59.28 77.10 ± 1.34 88.50 84.55 ± 0.06
Coll. foam 64.45 61.79 ± 2.14 68.22 88.08 ± 0.07
Coll. cables 52.17 59.55 ± 2.80 66.00 82.86 ± 0.19
Coll. cardboard 55.44 56.18 ± 4.64 68.87 91.17 ± 0.14
Var. can weight 53.71 69.28 ± 1.95 82.66 80.29 ± 0.10
Cable at robot 34.99 76.52 ± 5.25 100.00 98.89 ± 0.04
Invalid grip. pos. 58.63 74.71 ± 6.05 90.97 89.29 ± 0.14
Unstable platform 60.80 73.23 ± 2.51 79.97 94.65 ± 0.08
Mean ± SD 57.23 ± 8.48 73.66 ± 12.05 84.27 ± 11.55 89.23 ± 7.45
The mean represents the average score across all anomaly types, and SD stands for the standard deviation among all anomaly types. The variances of the results are 0.0 for Anomaly transformer and MVT-Flow because they are deterministic. Consequently, only the mean values are reported for these models.
Table 4: AUROC percentages of real-time anomaly detection models on 50 Hz torque signals trained with 948 (100%) samples.
Model
Anomaly
transformer
LSTM-VAE
MVT-Flow
(previous SOTA)
Sparse MAF-AAE
Add. Friction 68.81 83.99 ± 1.03 80.99 93.24 ± 0.08
Miscommutation 58.87 69.38 ± 1.99 83.90 74.86 ± 0.24
Misgrip can 68.37 97.92 ± 0.70 98.00 99.83 ± 0.02
Losing can 52.78 91.75 ± 1.04 82.23 89.04 ± 0.19
Add. axis weight 65.81 73.37 ± 1.09 79.82 82.39 ± 0.15
Coll. foam 64.44 69.22 ± 1.71 68.16 77.18 ± 0.16
Coll. cables 51.68 61.74 ± 2.29 52.51 72.89 ± 0.26
Coll. cardboard 58.19 70.44 ± 2.78 57.26 91.44 ± 0.33
Var. can weight 59.32 83.16 ± 1.32 68.68 84.75 ± 0.21
Cable at robot 62.94 83.96 ± 3.28 100.00 93.79 ± 0.22
Invalid grip. pos. 62.05 72.16 ± 3.27 86.22 95.64 ± 0.18
Unstable platform 59.25 69.16 ± 1.70 70.85 90.66 ± 0.19
Mean ± SD 61.04 ± 5.46 77.19 ± 10.80 77.39 ± 14.54 87.14 ± 8.69
The mean represents the average score across all anomaly types, and SD stands for the standard deviation among all anomaly types. For the same reason as in Table  3, only the mean values are reported for Anomaly transformer and MVT-Flow.
For a detailed comparison, we calculated the AUROC percentages of each anomaly type and the average performance of each model on 100 Hz and 50 Hz torque signals, respectively, trained with 948 (100%) samples. Figure 5 and Figure 6 showcase the ROC curves for each initial run on 100 Hz and 50 Hz, respectively. The numerical results are presented in Table 3 and 4. Sparse MAF-AAE outperformed others in 7 out of 12 categories at 100 Hz and 9 out of 12 at 50 Hz. Furthermore, our model showed the smallest variance across all anomalies when compared to LSTM-VAE and MVT-Flow. This consistency indicates that our model not only excels in detecting a wide range of anomalies but also maintains superior and uniform performance across all categories, underscoring its robustness and reliability.

### v3

@@ Abstract
The growing demand for robots to operate effectively in diverse environments necessitates the need for robust real-time anomaly detection techniques during robotic operations. However, deep learning-based models in robotics face significant challenges due to limited training data and highly noisy signal features. In this paper, we present Sparse Masked Autoregressive Flow-based Adversarial AutoEncoder model to address these problems. This approach integrates Masked Autoregressive Flow model into Adversarial AutoEncoders to construct a flexible latent space and utilize Sparse autoencoder to efficiently focus on important features, even in scenarios with limited feature space. Our experiments demonstrate that the proposed model achieves a 4.96% to 9.75% higher area under the receiver operating characteristic curve for pick-and-place robotic operations with randomly placed cans, compared to existing state-of-the-art methods. Notably, it showed up to 19.67% better performance in scenarios involving collisions with lightweight objects. Additionally, unlike the existing state-of-the-art model, our model performs inferences within 1 millisecond, ensuring real-time anomaly detection. These capabilities make our model highly applicable to machine learning-based robotic safety systems in dynamic environments. The code is available at https://github.com/twkang43/sparse-maf-aae.
@@
journal: Engineering Applications on Artificial Intelligence
[…]
10.1016/j.engappai.2025.111310
@@ 1 Introduction
Neural network-based time-series anomaly detection approaches have been actively researched in recent times (Pang et al., 2021). Unsupervised learning, which uses only normal data while training the model, is predominantly used in this field (Pang et al., 2021) since the occurrence of anomalies is very rare and significant effort is required to identify and label data. Models such as Long Short-Term Memory-based Variational AutoEncoder (LSTM-VAE) (Park et al., 2018), MultiVariate Time-series Flow (MVT-Flow) (Brockmann et al., 2024), Anomaly transformer (Xu et al., 2022) are proposed to efficiently detect anomalies using unsupervised learning.
@@
Anomaly transformer achieves success in the field of anomaly detection by utilizing transformer encoders (Vaswani et al., 2017) to capture both temporal and global time-series features of the data. It can effectively detect anomalies because the transformer can capture global time-series information that recurrent neural networks are unable to retain. However, due to the algorithm’s time and space complexity of $O(L^{2}N)$, where $L$ represents the length of the time-series sequence and $N$ denotes the number of features, it requires substantial computational time and memory space. This becomes a drawback when performing real-time anomaly detection with long time steps. In robotics, the high computational cost of the method is a critical weakness since it has to be implemented on the embedded computing unit for real-time anomaly detection. Specifically, using numerous sensor signals increases the computational load, rendering it unsuitable for use on robots, particularly low-cost commercial robots (Yen et al., 2019). Additionally, the model’s large size requires extensive datasets for training, but building these datasets through robot operation is challenging due to the time-intensive data collection process. (Soori et al., 2023).
@@
In this paper, we propose a real-time multivariate time-series anomaly detection model, Sparse Masked Autoregressive Flow-based Adversarial AutoEncoder (Sparse MAF-AAE), designed to address the limitations of conventional anomaly detection models. First, we enhance the adaptability of the model’s latent space for datasets, designing it to have a flexible posterior and prior distributions. This flexibility is achieved by transforming the posterior and prior distributions within Adversarial AutoEncoder (AAE) (Makhzani et al., 2016) using Masked Autoregressive Flow (MAF) (Papamakarios et al., 2017), a type of NF. The combination of AAE and MAF enables the model to learn a latent space that captures intricate data structures, performing well even with a small amount of training data. Additionally, this approach enhances robustness against out-of-distribution data by integrating Masked Autoencoder for Distribution Estimation (MADE) (Germain et al., 2015) within MAF, utilizing a bottleneck architecture (Kirichenko et al., 2020). Second, to preserve essential signal characteristics when the number of signal features is small, we incorporate the concept of Sparse autoencoder (Ng et al., 2011) by creating a latent space larger than the input while applying constraints that limit parameter updates. This configuration enables logical compression of the input data and ensures effective feature preservation for constructing an appropriate latent space from sparse input sensor signals. As a result, it enhances the model’s effectiveness in detecting anomalies in robotic operations with limited sensor signals. Finally, we demonstrated that our methods are feasible for real-time multivariate anomaly detection in robots by constructing models with simple architecture, achieving suitable elapsed times for real-time inferences. To the best of our knowledge, this is the first work to integrate a flow-based model, AAE, and Sparse autoencoder within a unified framework for signal reconstruction-based anomaly detection.
@@
Comparative experiments are conducted to evaluate the real-time multivariate anomaly detection performance based on a modified version of the voraus-AD dataset (Brockmann et al., 2024). The dataset offers time-series signals of the manipulator when a pick-and-place operation is performed with 12 anomaly types caused by heterogeneous factors. To implement the real-time detection, the sliding window is applied for the implemented anomaly detection methods. Results showed that our model achieved the highest area under the ROC curve (AUROC) score among the comparison models, while also maintaining an appropriate computational time for real-time multivariate time-series anomaly detection in robotic applications.
@@ 4.3 Implementation Details
Table 3: AUROC percentages of online anomaly detection models on 100 Hz and 50 Hz torque signals
trained with 948 (100%) samples.
Frequency 100 Hz 50 Hz
Model
Anomaly
transformer
LSTM-VAE MVT-Flow
Sparse
MAF-AAE
Anomaly
transformer
LSTM-VAE MVT-Flow
Sparse
MAF-AAE
Add. Friction 69.63 81.60 ± 1.52 87.20 96.29 ± 0.03 68.81 83.99 ± 1.03 80.99 93.24 ± 0.08
Miscommutation 58.35 67.53 ± 1.75 90.96 75.87 ± 0.11 58.87 69.38 ± 1.99 83.90 74.86 ± 0.24
Misgrip can 62.53 97.99 ± 2.28 99.74 99.82 ± 0.02 68.37 97.92 ± 0.70 98.00 99.83 ± 0.02
Losing can 56.83 88.48 ± 1.35 88.20 89.03 ± 0.11 52.78 91.75 ± 1.04 82.23 89.04 ± 0.19
Add. axis weight 59.28 77.10 ± 1.34 88.50 84.55 ± 0.06 65.81 73.37 ± 1.09 79.82 82.39 ± 0.15
Coll. foam 64.45 61.79 ± 2.14 68.22 88.08 ± 0.07 64.44 69.22 ± 1.71 68.16 77.18 ± 0.16
Coll. cables 52.17 59.55 ± 2.80 66.00 82.86 ± 0.19 51.68 61.74 ± 2.29 52.51 72.89 ± 0.26
Coll. cardboard 55.44 56.18 ± 4.64 68.87 91.17 ± 0.14 58.19 70.44 ± 2.78 57.26 91.44 ± 0.33
Var. can weight 53.71 69.28 ± 1.95 82.66 80.29 ± 0.10 59.32 83.16 ± 1.32 68.68 84.75 ± 0.21
Cable at robot 34.99 76.52 ± 5.25 100.00 98.89 ± 0.04 62.94 83.96 ± 3.28 100.00 93.79 ± 0.22
Invalid grip. pos. 58.63 74.71 ± 6.05 90.97 89.29 ± 0.14 62.05 72.16 ± 3.27 86.22 95.64 ± 0.18
Unstable platform 60.80 73.23 ± 2.51 79.97 94.65 ± 0.08 59.25 69.16 ± 1.70 70.85 90.66 ± 0.19
Mean ± SD 57.23 ± 8.48 73.66 ± 12.05 84.27 ± 11.55 89.23 ± 7.45 61.04 ± 5.46 77.19 ± 10.80 77.39 ± 14.54 87.14 ± 8.69
The mean represents the average score across all anomaly types, and SD stands for the standard deviation among all anomaly types.
The variance for each type of result produced by the Anomaly transformer and MVT-Flow models is 0.0, as these models are deterministic in nature.
Consequently, only the mean values are reported for each type in these models.
@@ 4.4 Results
For a detailed comparison, we calculated the AUROC percentages of each anomaly type and the average performance of each model on 100 Hz and 50 Hz torque signals, respectively, trained with 948 (100%) samples. Figure 5 and Figure 6 showcase the ROC curves for each initial run on 100 Hz and 50 Hz, respectively. The numerical results are presented in Table 3. Sparse MAF-AAE outperformed others in 7 out of 12 categories at 100 Hz and 9 out of 12 at 50 Hz. Furthermore, our model showed the smallest variance across all anomalies when compared to LSTM-VAE and MVT-Flow. This consistency indicates that our model not only excels in detecting a wide range of anomalies but also maintains superior and uniform performance across all categories, underscoring its robustness and reliability.